# ESVO2: Direct Visual-Inertial Odometry with Stereo Event Cameras

### v1

@@ Abstract
Event-based visual odometry is a specific branch of visual Simultaneous Localization and Mapping (SLAM) techniques, which aims at solving tracking and mapping sub-problems in parallel by exploiting the special working principles of neuromorphic (i.e., event-based) cameras. Due to the motion-dependent nature of event data, explicit data association (i.e., feature matching) under large-baseline view-point changes is hardly established, making direct methods a more rational choice. However, state-of-the-art direct methods are limited by the high computational complexity of the mapping sub-problem and the degeneracy of camera pose tracking in certain degrees of freedom (DoF) in rotation. In this paper, we resolve these issues by building an event-based stereo visual-inertial odometry system on top of our previous direct pipeline Event-based Stereo Visual Odometry [1]. Specifically, to speed up the mapping operation, we propose an efficient strategy for sampling contour points according to the local dynamics of events. The mapping performance is also improved in terms of structure completeness and local smoothness by merging the temporal stereo and static stereo results. To circumvent the degeneracy of camera pose tracking in recovering the pitch and yaw components of general six-DoF motion, we introduce IMU measurements as motion priors via pre-integration. To this end, a compact back-end is proposed for continuously updating the IMU bias and predicting the linear velocity, enabling an accurate motion prediction for camera pose tracking. The resulting system scales well with modern high-resolution event cameras and leads to better global positioning accuracy in large-scale outdoor environments. Extensive evaluations on five publicly available datasets featuring different resolutions and scenarios justify the superior performance of the proposed system against five state-of-the-art methods. We release our pipeline as an open-source software for future research in this field.
@@
Supplemental video: https://youtu.be/mdhomp3OCbE
Code: https://github.com/NAIL-HNU/ESVO2.git
@@ I Introduction
Drawing inspiration from biological systems, event cameras represent a revolutionary advancement in sensor technology. These sensors transmit pixel-wise intensity changes, called “events”, asynchronously as they occur [3, 4]. Consequently, they do not produce grayscale images, nor do they adhere to an external clock of fixed capture rate typical of traditional cameras. This asynchronous and differential principle of operation minimizes temporal redundancy, thereby reducing power consumption and bandwidth requirements. Endowed with high temporal resolution and high dynamic range (HDR) capabilities [4], event cameras are qualified to deal with challenging scenarios that are inaccessible to traditional cameras, such as high-speed motion and/or HDR illumination tracking [5, 6, 7, 8, 9, 10, 11, 12], control [13, 14, 15], image synthesis [16, 17, 18, 19] and Simultaneous Localization and Mapping (SLAM) [20, 21, 22, 23, 24, 25, 26, 27].
Although the potential of event cameras is promising, their unconventional output is incompatible with existing modules in visual odometry (VO) / SLAM systems designed for traditional cameras. Event-based VO, on the one hand, inherits the theoretical formulation and parallel design from modern VO solutions. On the other hand, the distinctiveness of event-based methods is witnessed in several aspects, including data processing and representation. Early works using a monocular event camera (i.e., [21, 23]) require a very gentle motion (typically a local-loopy behavior) for the initialization of a local 3D map, based on which the camera pose can be tracked using a 3D-2D registration pipeline. The dependence on dedicated hardware (e.g., a GPU) [21] and the inefficient map expansion [21, 23] restrict their applications on mobile platforms equipped with a limited computation resource and energy supply. More recently, Zhou et al. [1] have proposed an event-based stereo visual odometry (ESVO) pipeline that achieves better accuracy and demonstrates a real-time performance on data collected by a moving agent, such as a flying drone or a ground vehicle. However, this method does not scale well with the streaming rate of event cameras. Real-time performance is only witnessed on event cameras with a small spatial resolution (Q-VGA), e.g., a DAVIS346 (Dynamic and Active-Pixel Vision Sensor) [28]. Also, the failure of tracking in the presence of a sudden variation in angular velocity casts the necessity of introducing an inertial measurement unit (IMU) into the system.
@@ II Related Work
We improve the AA scheme by incorporating the approach from [56], allowing it to adapt to the variation of event rate without manually adjusting parameters.
@@
We significantly reduce the computational cost in the stereo matching by accelerating the stereo similarity evaluation, enabling the mapping module to run stably at 20 Hz on event data of VGA spatial resolution.
@@
We introduce a back-end using as input the IMU data and past poses to estimate IMU biases and the camera’s velocity, which are then used to provide initial values for the spatio-temporal registration, leading to more efficient and accurate performance of camera pose tracking.
@@
The proposed pipeline is more extensively evaluated, and the performance in terms of efficiency and accuracy is further improved.
@@ IV-A Pre-processing: Adaptive Accumulation of Events
The method in [56] utilizes the event activity within a time interval to represent the global event dynamics, controlling the decay rate depending on the event activity to generate adaptive-decay time surfaces. Inspired by [56], our method monitors the event activity to determine the termination time of event accumulation. Intuitively, a greater event activity corresponds to higher event dynamics, indicating a shorter duration for event accumulation. We also observe that the event dynamics remain relatively stable within a short time interval (e.g., 10 ms), which leads to event activities converging towards a constant value incrementally, as shown in Fig. 0(a). Assuming the events are generated at equal time intervals (e.g., s), the converged event activity can be roughly calculated as
As seen in Fig. 0(b), this estimation of converged event activity also holds approximately when the time interval is not constant. Therefore, we terminate event accumulation when the event activity converges to obtain a sufficiently sharp image. To handle the spatial variation of local event dynamics, we divide the image plane into small regions and accumulate events independently according to the local event activity.
The generation of an AA map is summarized in Algorithm 1. Specifically, we first divide the image plane evenly into small regions, and the accumulation of events is carried out in each region independently. For each region, the converged event activity is estimated using (3), and then all spatially involved events are sequentially accumulated and used to update the event activity until it converges. The pixel value in an AA map represents the number of events accumulated at this pixel. Pixels with higher values are more likely to be associated with contour points.
To assess the effectiveness of the proposed AA method, we compare it against three image-like representations, including time surfaces [58] and two speed-invariant ones, namely Speed Invariant Learned Corners (SILC) [59] and Threshold-Ordinal Surface (TOS) [60]. As illustrated in Fig. 1, the AA map preserves relatively complete edges with fewest redundant points while keeping the highest signal-to-noise ratio (SNR)222By “SNR”, we mean the ratio between edge points and non-edge points in all the sampled points.. This superior performance can also be justified quantitatively by comparing each representation against the edge map obtained from the corresponding gray-scale image. As shown in Tab. II, the proposed AA representation leads to the highest score using the multi-component structure similarity index measure (3-SSIM) [61] as evaluation metric.
@@
Contour points used by the following mapping operation are further sampled from each region independently. In general, the higher the AA pixel value in each region, the greater the probability that the pixel will be selected. We compare the sampling result with contour points obtained from ESVO. As shown in Fig. 2, our sampled pixels better capture the edge structures, indicating that our sampling method can cover the edge patterns in the environment using fewer redundant input points.
@@ IV-B Mapping: Depth Estimation
The mapping module aims at recovering depth information and maintaining a local 3D map. In the original ESVO pipeline, these are achieved by constantly propagating recent depth estimates of sampled events to a virtual reference frame, on which overlapped estimates are further fused based on their probabilistic characteristics. As for estimating the depth of each individual event, a two-step method (i.e., a block matching operation followed by a nonlinear refinement) is proposed. In general, the applied static stereo configuration leads to unreliable depth estimation of structures parallel to the stereo baseline. Moreover, the two-step method for depth estimation does not scale well with the number of individual queries. To circumvent these two problems, we propose two modifications: 1) an additional temporal-stereo matching method and 2) a faster design of block matching. We detail each modification in the following.
@@ IV-B1 Temporal Stereo Matching
Let’s consider the commonly used horizontal stereo configuration. We divide the sampled contour points from Sec. IV-A into two groups according to their gradient direction on the TS. In the first group, we collect pixels at which the ratio () between the vertical gradient and the horizontal gradient is smaller than a threshold. These pixels are fed to the static stereo method. The remaining sampled pixels go to the second group and are fed to the proposed temporal stereo method.
The key to the event-based temporal-stereo problem is to effectively exploit appearance similarity in event data. This requires the applied event representation, on which the stereo data association is established, to possess the speed-invariant property to some extent. Therefore, the TS used in ESVO [1] is not applicable. To address this, we investigate the temporal stereo matching performance on three representations, namely TOS [60], SILC [59], and our AA. We find that AA is the optimal choice because of its highest contour points ratio. Consequently, the epipolar matching is carried out on successive AA maps of the left camera.
@@ IV-B2 Fast Block Matching
| Average time | 0.0317 ms | 0.0057 ms |
Although the original ZNCC leads to satisfactory accuracy, it is notably time-consuming. To speed up, we propose a fast version of ZNCC based on the approach presented in [64]. It adopts an recursive method that calculates the required sums and squared sums for the computation of ZNCC:
@@ IV-B3 Depth Merge
where , , are the model parameters, namely the mean, scale, and degree of freedom. However, this is not sensible because the uncertainty of the temporal-stereo results is always much smaller, leading to a bias towards the temporal-stereo estimates. Hence, we simply merge the results of the two stereo methods to obtain a more complete depth map, as shown in Sec. VI-B. Also note that we get rid of the nonlinear refinement of depth in the static-stereo and temporal-stereo methods in our implementation. We find that the absence of the nonlinear refinement guarantees the real-time performance of mapping, while not bringing any notable degradation in the accuracy of camera pose tracking. The effect of omitting the nonlinear refinement is discussed in Sec. VI-E3.
@@ IV-C Localization: Camera Pose Estimation
Our front-end basically follows the tracking module in ESVO [1], which takes as input a TS and a local 3D map, and computes the pose of the stereo rig with respect to the map. Let represent a set of pixels with inverse depth values in the reference frame, and and be the TS and negative TS at time . The objective is, identical to [1], to determine the optimal motion parameters by solving
where the warp function denotes the transformation from local depth points to the latest negative TS image. Besides, denotes the motion parameters, where are the Cayley parameters [65] for rotation and represents the translation.
@@ IV-C2 Using IMU Pre-integration as Motion Priors
where , , and are the bias and noise of the accelerometer and gyroscope in the IMU frame, respectively. The relative translation, velocity, and rotation between two successive pose estimates, e.g., from time to , expressed in the IMU frame , can be calculated by
where denotes the gravity in the IMU coordinate system at time , , and the skew-symmetric matrix of a vector. The linear velocity and biases (i.e., , and ) are initialized empirically and constantly updated in the back-end (see V). The relative velocity and rotation from IMU frame to , namely and , are used as motion priors in the spatio-temporal registration.
@@ V Back-end Optimization
To suppress the drift in the motion estimates, a bundle adjustment (BA) optimization is typically needed as a back-end. One of the most widely used strategies is the sliding-window approach, which jointly refines the poses of all involved key frames and, optionally, the depth information of co-visible 3D points by minimizing the inconsistency among heterogeneous measurements. However, our geometric method lacks explicit data association in the spatio-temporal domain, and thus, such a relaxation-and-refinement scheme [67] is not applicable. Moreover, we observe that the presence of camera poses among the optimization variables does not improve the camera tracking performance. This is because the local map (i.e., the involved point cloud within the window) is a fusion result relying on existing camera poses. It is by no means to establish additional constraints by introducing the 3D point cloud into the sliding-window optimization without relaxation. Therefore, we attempt to create a compact and efficient back-end, which considers only the linear velocity and IMU’s bias as the optimizing variables. Experiments show that the back-end built on this strategy significantly improves the tracking accuracy of our system. (Sec VI-E)
Specifically, we apply a sliding window (shown in Fig. 4), in which the last pose estimates333Note that the involved poses are only used for providing the relative pose constraints rather than being updated. and the corresponding IMU’s measurements are involved. The objective is to refine the estimated linear velocity and IMU biases at the corresponding time of each involved pose so that the relative poses comply maximally with the IMU pre-integration results:
where denotes the optimizing variables (i.e., the linear velocities and the biases of IMU), the involved pose estimates, and the IMU pre-integrated measurements. In addition, denotes the set of IMU pre-integration results in the sliding window. The relative pose residual is specifically defined as:
where and denote the position and rotation, respectively. extracts the vector portion of the quaternion for representing the error state. In detail, the size of the sliding window is set to 5, and we apply the Levenberg-Marquardt (LM) algorithm to solve the optimization problem. The sliding window shifts forward when a new pose estimate is obtained, while the oldest pose estimate and corresponding IMU measurements are removed.

### v3

@@ Abstract
Event-based visual odometry is a specific branch of visual Simultaneous Localization and Mapping (SLAM) techniques, which aims at solving tracking and mapping sub-problems (typically in parallel), by exploiting the special working principles of neuromorphic (i.e., event-based) cameras. Due to the motion-dependent nature of event data, explicit data association (i.e., feature matching) under large-baseline viewpoint changes is difficult to establish, making direct methods a more rational choice. However, state-of-the-art direct methods are limited by the high computational complexity of the mapping sub-problem and the degeneracy of camera pose tracking in certain degrees of freedom (DoF) in rotation. In this paper, we tackle these issues by building an event-based stereo visual-inertial odometry system on top of a direct pipeline [1]. Specifically, to speed up the mapping operation, we propose an efficient strategy for sampling contour points according to the local dynamics of events. The mapping performance is also improved in terms of structure completeness and local smoothness by merging the temporal stereo and static stereo results. To circumvent the degeneracy of camera pose tracking in recovering the pitch and yaw components of general 6-DoF motion, we introduce IMU measurements as motion priors via pre-integration. To this end, a compact back-end is proposed for continuously updating the IMU bias and predicting the linear velocity, enabling an accurate motion prediction for camera pose tracking. The resulting system scales well with modern high-resolution event cameras and leads to better global positioning accuracy in large-scale outdoor environments. Extensive evaluations on five publicly available datasets featuring different resolutions and scenarios justify the superior performance of the proposed system against five state-of-the-art methods. Compared to ESVO [1], our new pipeline significantly reduces the camera pose tracking error by – and –- in terms of absolute trajectory error and relative pose error, respectively; at the same time, the mapping efficiency is improved by a factor of five. We release our pipeline as an open-source software for future research in this field.
@@ I Introduction
Drawing inspiration from biological systems, event cameras represent a revolutionary advancement in sensor technology. These sensors transmit pixel-wise intensity changes, called “events”, asynchronously as they occur [3, 4, 5]. Consequently, they do not produce grayscale images, nor do they adhere to an external clock of fixed capture rate typical of traditional cameras. This asynchronous and differential principle of operation minimizes temporal redundancy, thereby reducing power consumption and bandwidth requirements. Endowed with high temporal resolution and high dynamic range (HDR) capabilities [4, 5], event cameras are qualified to deal with challenging scenarios that are inaccessible to traditional cameras, such as high-speed motion and/or HDR illumination tracking [6, 7, 8, 9, 10, 11, 12, 13], control [14, 15, 16, 17, 18, 19, 20, 21], image synthesis [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and Simultaneous Localization and Mapping (SLAM) [32, 33, 34, 35, 36, 37, 38, 39, 40].
Although the potential of event cameras is promising, their unconventional output is incompatible with existing modules in visual odometry (VO) / SLAM systems designed for traditional cameras. Event-based VO, on the one hand, inherits the theoretical formulation and parallel design from modern VO solutions. On the other hand, the distinctiveness of event-based methods is witnessed in several aspects, including data processing and representation. Early works using a monocular event camera (i.e., [33, 35]) require a very gentle motion (typically a small-loopy behavior) for the initialization of a local 3D map, based on which the camera pose can be tracked using a 3D-2D registration pipeline. The dependence on dedicated hardware (e.g., a GPU) [33] and the inefficient map expansion [33, 35] restrict their applications on mobile platforms equipped with a limited computation resource and energy supply. More recently, Zhou et al. [1] have proposed an event-based stereo visual odometry (ESVO) pipeline that achieves better accuracy and demonstrates a real-time performance on data collected by a moving agent, such as a flying drone or a ground vehicle. However, this method does not scale well with the streaming rate of event cameras. Real-time performance is only witnessed on event cameras with a small spatial resolution (Q-VGA), e.g., a DAVIS346 (Dynamic and Active-Pixel Vision Sensor) [41]. Also, the failure of tracking in the presence of a sudden variation in angular velocity casts the necessity of introducing an inertial measurement unit (IMU) into the system.
@@
This work has evolved from our ICRA paper [42], and the technical improvements are clarified at the end of Sec. II.
@@ II Related Work
The AA scheme originally proposed in [42] is non-trivially improved here. Compared to the original design, we introduce the idea of event-dynamic monitoring [69] into the termination criteria of event accumulation. This modification makes AA responsive to event rate variations without manually adjusting parameters. To achieve real-time performance, we also derive an approximate formula to calculate the converged event activity.
@@
Inheriting the combination of temporal-stereo and static-stereo configurations from [42], we significantly reduce the computational cost of static-stereo operation in this paper by proposing a fast block-matching scheme, which ensures the mapping module to run stably at 20 Hz on event data of VGA spatial resolution.
@@
Using IMU pre-integration as motion priors was naively proposed in [42], however, the IMU bias was not effectively suppressed. To solve this problem while ensuring real-time performance, we create a compact and efficient back-end, which considers only linear velocity and IMU biases as optimizing variables. Using the back-end’s results as the initial values for the spatio-temporal registration leads to more efficient and accurate performance of camera pose tracking. According to our investigation on various back-end designs (e.g., [63]), our model strikes a good balance between accuracy and efficiency.
@@
The proposed pipeline is more extensively evaluated. Compared to [42], ESVO2 reduces the trajectory error by approximately 50% across all datasets and improves the computational efficiency of the mapping subproblem by more than a factor of two.
@@ IV-A Pre-processing: Adaptive Accumulation of Events
The method in [69] utilizes the event activity within a time interval to represent the global event dynamics, controlling the decay rate depending on the event activity to generate adaptive-decay time surfaces. Inspired by [69], our method monitors the event activity to determine the termination time of event accumulation. Intuitively, a greater event activity corresponds to higher event dynamics, indicating a shorter duration for event accumulation. We also observe that the event dynamics remain relatively stable within a short time interval (e.g., 10 ms), which leads to event activities converging towards a constant value incrementally, as shown in Fig. 0(a). Assuming the events are generated at equal time intervals, the converged event activity can be roughly calculated as
where denotes the length of the time interval in seconds. As seen in Fig. 0(b), this estimation of converged event activity also holds approximately when the time interval is not constant. Therefore, we terminate event accumulation when the event activity converges to obtain a sufficiently sharp image. To handle the spatial variation of local event dynamics, we divide the image plane into small regions and accumulate events independently according to the local event activity.
The generation of an AA map is summarized in Algorithm 1. Specifically, we first divide the image plane evenly into small regions and the accumulation of events is carried out in each region independently. Events are sequentially accumulated in a region of an event count image until the event activity converges to the steady state value estimated by (3). The pixel value in an AA map represents the number of events accumulated at this pixel. Pixels with higher values are more likely to be associated with contour points.
To assess the effectiveness of the proposed AA method, we compare it against three image-like representations, including time surfaces [71] and two speed-invariant ones, namely Speed Invariant Learned Corners (SILC) [72] and Threshold-Ordinal Surface (TOS) [73]. As illustrated in Fig. 1, the AA map preserves relatively complete edges with fewest redundant points while keeping the highest true-edge-point ratio (TEPR) 222By “TEPR”, we mean the proportion of contour points covered by the sampled points, as illustrated in Fig. 2.. This superior performance can also be justified quantitatively by comparing each representation against the edge map obtained from the corresponding grayscale image. As shown in Tab. II, the proposed AA representation leads to the highest score using the multi-component structure similarity index measure (3-SSIM) [74] as evaluation metric.
@@
Contour points are further sampled from each region independently. In general, the higher the AA pixel value in each region, the greater the probability that the pixel will be selected. We compare the sampling result to the contour points obtained by ESVO. As shown in Fig. 2, our sampled pixels better capture the edge structures, indicating that our sampling method can cover the edge patterns in the environment using fewer redundant input points.
@@ IV-B Mapping: Depth Estimation
The mapping module aims at recovering depth information and maintaining a local 3D map. In the original ESVO pipeline, these are achieved by constantly propagating recent depth estimates of sampled events to a virtual reference frame, on which overlapped estimates are further fused based on their probabilistic characteristics. As for estimating the depth of each individual event, a two-step method (i.e., a block matching operation followed by a nonlinear refinement) is proposed. More specifically, a triangulation operation is performed given the resulting disparity from the block matching, and the nonlinear refinement operates in a forward-projection manner (i.e., from the inverse depth domain to 2D event measurements). In general, the applied static stereo configuration in ESVO leads to unreliable depth estimation of structures parallel to the stereo baseline. Moreover, the two-step method for depth estimation does not scale well with the number of individual queries. To circumvent these two problems, we propose two modifications: 1) an additional temporal-stereo matching method and 2) a faster design of block matching. We detail each modification in the following.
@@ IV-B1 Temporal Stereo Matching
Let us consider the commonly used horizontal stereo configuration. We divide the sampled contour points from Sec. IV-A into two groups according to their gradient direction on the TS. In the first group, we collect pixels at which the ratio () between the vertical gradient and the horizontal gradient is smaller than a threshold. These pixels are fed to the static stereo method. The remaining sampled pixels go to the second group and are fed to the proposed temporal stereo method.
The key to the event-based temporal-stereo problem is to effectively exploit appearance similarity in event data. This requires the applied event representation, on which the stereo data association is established, to possess the speed-invariant property to some extent. Since the TS used in ESVO [1] does not satisfy this property, it is not a good choice for temporal stereo matching. To address this, we investigate the temporal stereo matching performance on three representations, namely TOS [73], SILC [72], and our AA. We find that AA is the optimal choice due to its highest contour points ratio. Consequently, the epipolar matching is carried out on successive AA maps of the left camera, followed by triangulation and optionally by nonlinear refinement.
Note that the complementary property of the temporal-stereo configuration to the static one would hold as long as the temporal-stereo epipolar line is not parallel to the static-stereo baseline. This assumption is typically valid for most mobile robot platforms (e.g., ground vehicles and drones) when the stereo camera is set looking forward.
@@ IV-B2 Fast Block Matching
| Mean time | 30.70 s | 5.761 s |
| Median time | 30.16 s | 5.714 s |
| 25th percentile | 29.97 s | 5.571 s |
| 75th percentile | 30.45 s | 5.819 s |
| Standard deviation | 2.369 s | 0.452 s |
Although the original ZNCC leads to satisfactory accuracy, it is notably time consuming. To speed up, we propose a fast version of ZNCC based on the approach presented in [77]. It adopts a recursive method that calculates the required sums and squared sums for the computation of ZNCC:
@@ IV-B3 Depth Merge
where , , are the model parameters, namely the mean, scale, and degree of freedom. However, this is not sensible because the uncertainty of the temporal-stereo results is always much smaller, leading to a bias towards the temporal-stereo estimates. Hence, we straightforwardly merge the results of the two stereo methods to obtain a more complete depth map, as shown in Sec. VI-B. Specifically, the employed merging operation consists of two steps. First, the depth estimates from the static-stereo and temporal-stereo methods are fused independently into two separate depth maps according to their respective probabilistic characteristics, in the same way as described in [1]. Afterwards, we fill the non-occupied pixels in the static-stereo depth map with the results in the temporal-stereo depth map. Also note that we get rid of the nonlinear refinement of depth in the static-stereo and temporal-stereo methods in our implementation. We find that the absence of the nonlinear refinement guarantees the real-time performance of mapping, while not bringing any notable degradation in the accuracy of camera pose tracking. The effect of omitting the nonlinear refinement is discussed in Sec. VI-E3.
@@ IV-C Localization: Camera Pose Estimation
Our front-end follows the tracking module in ESVO [1], which takes as input a TS and a local 3D map, and computes the pose of the stereo rig with respect to the map. Let represent a set of pixels with inverse depth values in the reference frame, and and be the TS and negative TS at time . The objective is, as in [1], to determine the optimal motion parameters by solving
where the warp transfers points from the reference frame to the latest TS using the inverse depth and projection operator (see [1]). The transformation matrix corresponds to the motion parameters , where are the Cayley parameters [78] for rotation and are translation parameters.
@@ IV-C2 Using IMU Pre-integration as Motion Priors
where , , and are the bias and noise of the accelerometer and gyroscope in the IMU frame, respectively. The IMU measurement, e.g., from time to , expressed in the IMU frame , can be calculated by
where , and is the cross-product matrix operator of a vector. The linear velocity and biases (i.e., , and ) are initialized empirically and constantly updated in the back-end (see Sec. V). We estimate the position prior from to by , while neglecting the impact of acceleration due to the high operating rate of our tracking module, and employ as rotation prior. The prior position and orientation are used to initialize the search for the optimal pose in (10).
@@ V Back-end Optimization
To suppress the drift in the motion estimates, a bundle adjustment (BA) optimization is typically needed as a back-end. One of the most widely used strategies is the sliding-window approach, which jointly refines the poses of all involved key frames and, optionally, the depth information of co-visible 3D points by minimizing the inconsistency among heterogeneous measurements. However, our geometric method lacks explicit data association in the spatio-temporal domain, and thus, such a relaxation-and-refinement scheme [80] is not applicable. Moreover, we observe that the presence of camera poses among the optimization variables does not improve the camera tracking performance. This is because the local map (i.e., the involved point cloud within the window) is a fusion result relying on existing camera poses. It is by no means to establish additional constraints by introducing the 3D point cloud into the sliding-window optimization without relaxation. Therefore, unlike [81, 63], we attempt to create a compact and efficient back-end, which considers only the linear velocity and IMU’s bias as the optimizing variables. Experiments show that the back-end built on this strategy significantly improves the tracking accuracy of our system (Sec VI-E).
Specifically, we apply a sliding window (shown in Fig. 4), in which the last pose estimates333Note that the involved poses are only used for providing the relative pose constraints rather than being updated. and the corresponding IMU’s measurements are involved. The objective is to refine the estimated linear velocity and IMU biases at the corresponding time of each involved pose by minimizing the Mahalanobis norm of IMU measurement residuals:
where denotes the optimizing variables (i.e., the linear velocities and the biases of IMU), the involved pose estimates, and the IMU pre-integrated measurements. In addition, denotes the set of IMU pre-integration results in the sliding window. Furthermore, is the covariance matrix of the IMU pre-integration during the propagation process, and it is used as the weight for the IMU measurements. Detailed derivation can be found in Sec. IV.B of [81]. The relative pose residual is specifically defined as:
where indicates the gravity in the world coordinate system, and denote the position and rotation, respectively. extracts the vector portion of the quaternion for representing the error state. In detail, the size of the sliding window is set to 5, and we apply the Levenberg-Marquardt (LM) algorithm to solve the optimization problem. The sliding window shifts forward when a new pose estimate is obtained, while the oldest pose estimate and corresponding IMU measurements are removed.